OpenStudy (anonymous):

Justify the first step Akeem took in solving the equation below. 4(2x + 5) = –4 8x + 20 = –4 A. Akeem multipled both sides by 4. B. Akeem used the distributive property to eliminate the parentheses. C. Akeem subtracted 5 from both sides. D. Akeem added 4 to both sides.

OpenStudy (anonymous):

The best thing to do in this case is to replicate what was done in the first step. What did you do to get 8x + 20 = –4 ?

OpenStudy (anonymous):

If you're not sure, take a look at how -4 on the right side stays the same. So that must mean that whatever step was done was only done to the left side. That eliminates several of the options. What do you have left?

OpenStudy (mathmale):

In other words: 4(2x + 5) = –4 became 8x + 20 = –4 ... after doing what, and what was the reason for doing this?